interpret the slope
a.) y=2/3x+5
b.) y=-20x/9+4

Answer by stanbon(75887)

interpret the slope
a.) y=2/3x+5
slope = 2/3
As x increases 1, y increases by 2/3.
=====================================
b.) y=-20x/9+4
---
y = -(20/9)x + 4
slope = -20/9
As x increases 1, y decreases by -20/9
=====================================
Cheers,
Stan H.